OpenStudy (samsan9):

solve the equation for x: logx+log(x+2)=log(3x) A.) there is no solutions B.) x=1 C.) x=0 D.) x=0,1

OpenStudy (mathmate):

log(x)+log(x+2)=log(3x) log(x(x+2))=log(3x) => x(x+2)=3x (if x>0) Solve for x.

OpenStudy (anonymous):

logx+log(x+2)=log(3x) 1. Simplify logx+log(x+2) to logx(x+2) logx(x+2)=log3x 2. Cancel log on both sides x(x+2)=3x 3. Expand x^2+2x=3x 4. Move all terms to one side x^2+2x−3x=0 5. Simplify x^2+2x−3x to x2−x x^2−x=0 6. Factor out the common term x x(x−1) 7. Solve for x (Ask yourself, When will x(x−1) equal zero?)

OpenStudy (samsan9):

i think its when you have to equal x=o and x-1=0 so that it would be -1 or am i wrong?

OpenStudy (samsan9):

is the answer D?

OpenStudy (mathmate):

@samsan9 Recall that you have log(3x) on the right hand side, and the domain of log(x) is x>0. So when you solve the quadratic, be sure to reject all x<0 or x=0.

OpenStudy (samsan9):

so it would mean that x=1 since it seems like you are telling me that x shouldn't equal zero

OpenStudy (mathmate):

Yes, I AM. Recall I put "if x>0" to take care of this situation.

OpenStudy (samsan9):

so it is B?

OpenStudy (anonymous):

Yes @samsan9

OpenStudy (samsan9):

thanks you too :)